ttrin:

National Income (NI) is the sum of the incomes that all individuals in an economy earn in the forms of wages, interest, rents, and profits. It excludes government transfer payments and is calculated before any deductions are taken for income taxes. During the period 1994-2004, the NI of Australia grew about 5.2% per year, measured in 2003 U.S. dollars. In 1994, the NI was $4 billion. Assuming this rate of growth continues, what will the NI of Australia be (in billions) in the year 2015? Select one: a. $52.48 b. $1.28 c. $11.60 d. $33.88

RZzOo:

We set the equation Y= a (1+r)^t Where a is the starting amount and r is the rate r= 0.052 a=4 t=2015-1994 = 21 y= 4(1.052)^21 =11.59 =11.60